# Nova: Generative Language Models for Assembly Code with Hierarchical Attention and Contrastive Learning

### v5

@@ 1 Introduction
Large language models pre-trained on source code have brought improvement in various software development domains (Chen et al., 2022a; Liu et al., 2023; Chen et al., 2023; Le et al., 2022; Jiang et al., 2023; Xia et al., 2023). However, these LLMs are not designed for or trained with assembly corpus, not achieving their full potential on binary code analysis tasks such as binary code similarity (Wang et al., 2022; Xu et al., 2023a), malware detection (Su et al., 2024), and binary code decompilation (Tan et al., 2024; Armengol-Estapé et al., 2024; Hosseini & Dolan-Gavitt, 2022).
@@
  • We develop Nova, a generative foundation LLM with hierarchical attention and contrastive learning for assembly. Nova outperforms state-of-the-art on binary decompilation by up to 14.84 – 21.58% higher Pass@k and on binary similarity detection by up to 6.17% Recall@1.
@@ 2 Approach
Figure 2 presents the overall approach of Nova. We build Nova on top of foundation models for source code (Guo et al., 2024) to utilize their source code and natural language generation ability. We first collect large assembly corpora (Section 2.1). Section 2.2 describes Nova’s hierarchical attention design. With the collected assembly corpora, we then pretrain Nova with language modeling and contrastive learning objectives (Section 2.3). Then, we fine-tune Nova on two important downstream tasks, binary code decompilation, and binary code similarity detection (Sections 2.4 and 2.5), to prove Nova’s effectiveness and benefits to the binary research domain.
@@
Table 1: Statistics (number of source code and assembly functions) of the pre-training datasets.
Datasets Source O0 O1 O2 O3 Total
@@ 2.1 Data Collection
We build our assembly data sets on top of existing source code datasets: The-Stack (Li et al., 2023) and the AnghaBench (da Silva et al., 2021). We compile the source code into executables with different optimization levels (i.e., O0, O1, O2 and O3), strip the executables to remove debug information, and disassemble them into assembly code. We treat every assembly function as a separate data sample. The breakdown statistics are in Table 2.
@@ 2.2 Hierarchical Self-Attention
Nova uses hierarchical self-attention that is specially designed to learn the low-information-density semantics in the long sequence of assembly code. Specifically, Nova learns the assembly code in an hierarchical way by providing a modified attention mask. Different from standard token-level attentions (Vaswani et al., 2023; Radford & Narasimhan, 2018; Radford et al., 2019; Brown et al., 2020), our hierarchical self-attention contains three categories at different levels of granularity.
@@
To sum up, the hierarchical self-attention breaks the semantics of assembly code into three parts. The intra-instruction attention captures the instruction summary, and the preceding-instruction attention captures the context with the preceding instruction. The inter-instruction attention learns the long dependencies across instructions on top of the “[INST]” labels that contain the instruction summary. Figure 3 (c) also shows how Nova’s hierarchical attention for assembly code is compatible with the standard self-attention for text and source code. As existing LLMs have shown good performance on text and source code using the standard self-attention, we keep the standard causal attention mask within and between any chunks of text or source code in the input (the light grey part shown in Figure 3 (c)). The attention from text or source code to assembly code (and vice versa) is only allowed for the “[INST]” tokens as they are designed to contain the assembly instruction summaries.
@@ 2.3 Contrastive Learning
Let $e^{s}_{f}$ be the embedding of function $f$ in $s$ form ($s=-1$ for source code, and $s\in[0,1,2,3]$ for O0 to O3 optimized assembly). For simplicity, let $S=[-1,0,1,2,3]$ be the domain of $s$. We use the average of all the “[INST]” tokens’ embedding as the embedding of the whole assembly function, as each “[INST]” token is supposed to capture the semantics of that instruction by the design of hierarchical self-attention. Functionality CL optimizes Nova with the constraint:
@@
Such constraints can be trained by optimizing the embeddings of a batch of functions, each function in two different forms. For the example in Figure 4 (b), there are two forms (source code and O0 assembly) of four functions. Once Nova encodes the batch of source code and assembly functions, we calculate the distance matrix $\{D_{ij}\}_{f_{i},f_{j}\in F}=\{d(e^{s}_{f_{i}},e^{t}_{f_{j}})\}$, and minimize the loss:
@@
This objective minimizes the distance between embeddings for the same function, i.e., the diagonal in the distance matrix.
Optimization CL: LLMs can be confused if being asked to directly connect a source code function to its O3-optimized assembly, due to their dramatically different syntax. Such a huge gap can be filled by learning how the source code is transformed to O0, O1, O2 and eventually to O3 assembly, as the optimization levels are ordered.
Higher-level optimization applies a super-set of optimization rules compared to lower-level optimization. Nova learns such order with the optimization CL objective, encoding differently-optimized assembly code orderly. Optimization CL optimizes Nova with the constraint: the more optimizations applied, the larger the difference between embeddings of optimized and unoptimized code. For instance, Figure 4 (c) and (d) illustrate that for the same function “cmp”, the distance between source code and assembly increases when the optimization level increases. Formally, optimization CL minimizes the following loss:
@@ 2.5 Task 2: Binary Code Similarity Detection
A widely used setting is taking a query assembly of the function $f^{q}$ that is compiled with one optimization level (denoted by $s$), and a pool of candidate assembly of $K$ functions (notated by $f^{p}_{i},$ $1\leq i\leq K$) compiled with a different optimization level (denoted by $t\neq s$). There exists a unique candidate assembly coming from the same source code as the query ($\exists!1\leq i\leq K,$ $f^{p}_{i}=f^{q}$, called the positive candidate). Nova is fine-tuned to encode these binaries, so that the positive candidate has the highest similarity with the query assembly among the pool. The learning objective is as follows:
@@ 3 Experimental Setup
This section describes the setup of pre-training and fine-tuning of Nova, as well as the existing baselines we compare Nova with, and the evaluation metrics we used in the two downstream tasks. Appendix A.2 contains additional details such as training hyper-parameters.
@@ 3.1 Pre-Training
We use the data collected from AnghaBench and The-Stack for pre-training. We pre-train Nova starting from DeepSeek-Coder (Guo et al., 2024), and the hierarchical attention is applied on half of the attention heads to balance between its effectiveness and the existing knowledge in the standard attention layers (Justified in Appendix A.3). Nova is pre-trained with language modeling for one epoch, followed by contrastive learning objectives for another epoch.
@@ 3.2 Fine-Tuning for Binary Code Decompilation
Baselines: Nova is compared with existing SOTA LLM4Decompile (Tan et al., 2024), Meta LLMCompiler (Cummins et al., 2024), and other general code LLMs (CodeLlama (Rozière et al., 2023), StarCoder2 (Lozhkov et al., 2024), DeepSeekCoder (Guo et al., 2024), Qwen2.5-Coder (Hui et al., 2024), Granite Mishra et al. (2024), LongCoder (Guo et al., 2023), GPT-3.5-Turbo, and GPT-4o). LLM4Decompile trains DeepSeekCoder using the same AnghaBench corpus for binary decompilation. Meta LLMCompiler trains CodeLlama models using LLVM IRs, X86, and ARM assembly code to optimize and translate binary code.
Evaluation: GPT-3.5-Turbo and GPT-4o are prompted with three-shot examples, while all the other LLMs and Nova sample 20 decompilations per assembly code, using the temperature of 0.2 and top_p of 0.95 (Chen et al., 2021). The generated decompilations are executed against the test cases. Finally, Pass@1 and Pass@10 (Chen et al., 2021) are reported.
@@ 3.3 Fine-Tuning for Binary Code Similarity Detection
Baselines: Nova is compared with jTrans (Wang et al., 2022), DiEmph (Xu et al., 2023a) and CodeArt (Su et al., 2024). jTrans is a Transformer (Vaswani et al., 2023) encoder trained on binaries with masked token prediction and jump target prediction tasks. DiEmph uses an instruction deemphasis technique to prevent the model from learning instruction distribution biases introduced by compilers. CodeArt proposes a regularized attention mask for encoder models to capture instructional semantics and data dependencies.
Evaluation: We randomly sample $K$ source code functions from each project, compile them into binaries with O0 and O3 optimization flags, and disassemble them into X86-64 assemblies. BCSD techniques encode these assemblies into embeddings (Nova uses the average last-layer hidden states of all the “[INST]” tokens in an assembly as its embedding). Then each O0 assembly is used as the query to calculate their similarity with the $K$ O3 candidate assemblies (using cosine similarity). Metric Recall@1 is the ratio of queries for which the candidate from the same source code has the highest similarity among all the candidates.
@@ 4.1 Binary Code Decompilation
LLM4Decompile-1B 15.30 8.26 9.36 8.38 10.33 21.79 15.23 16.17 13.70 16.72
Nova-1B 37.53 21.71 22.68 18.75 25.17 49.38 34.84 36.95 32.03 38.30
LLM4Decompile-6B 29.97 19.05 20.46 18.32 21.95 40.40 27.75 28.85 28.51 31.38
Nova-6B 48.78 30.58 30.85 27.23 34.36 57.47 47.45 43.03 39.68 46.91
@@
With the same model size, Nova-1B outperforms LLM4Decompile-1B with a 14.84% higher averaged Pass@1, and a 21.58% higher Pass@10. Nova-6B outperforms LLM4Decompile-6B with a 12.41% higher averaged Pass@1, and a 15.53% higher Pass@10. When compared with GPT-4o, an order of magnitude larger model, Nova-1B produces an 8.38% higher Pass@1 and 13.80% higher Pass@10. Examples of Nova’s correct decompilation are provided in Appendix A.4.
Comparison with Techniques Handling Long Input: Nova’s hierarchical attention design targets to address the low information density and long input challenge of assembly code. There are other techniques that handles long input challenges in text and source code, with Granite-3B-Code-Base-128K and LongCoder being the most related ones. Granite trains LLM on repository-level long inputs, which is an orthogonal approach with Nova’s approach (hierarchical attention and contrastive learning). We train Granite-3B-Code-128K with Nova’s approach, and Table 3 shows that Nova’s approach brings improvement to Granite over standard fine-tuning even if it has already been trained with long code data.
@@
Granite + Nova’s Fine-Tuning 31.04 14.57 14.70 13.66 18.49 39.57 21.23 21.77 19.82 25.60
@@
DeepSeekCoder + LongCoder’s Attention 34.59 19.07 19.72 17.34 22.68 42.19 32.37 32.86 29.04 34.12
DeepSeekCoder + Nova’s Attention 37.53 21.71 22.68 18.75 25.17 49.38 34.84 36.95 32.03 38.30
LongCoder combines window attention and global attention to learn long code input. We compare LongCoder’s attention design with Nova’s hierarchical attention design using DeepSeekCoder-1.3B as the backbone. Table 4 shows that Nova’s hierarchical attention is more effective in learning assembly code. Nova’s attention design considers the instruction-level local semantics and dependencies between different instructions, which fits better than fix-sized window attention to assembly code.
Ablation Study: We conduct an ablation study by comparing Nova-1B with the following models to show the effectiveness of contrastive learning objectives and hierarchical attention design:
  • Nova-CL-HA: Removing contrastive learning and hierarchical self-attention. This is simply training DeepSeekCoder-1.3B on the assembly corpus using language modeling. This can be viewed as our reproduction (retrain) of LLM4Decompile-1B using the same amount of data.
@@
Table 5: Ablation study of Nova-1B on HumanEval-Decompile.
@@
Benchmarks jTrans DiEmph CodeArt Nova-1B Nova-6B
@@
Benchmarks jTrans DiEmph CodeArt Nova-1B Nova-6B
@@
Benchmarks jTrans DiEmph CodeArt Nova-1B Nova-6B
@@
Benchmarks jTrans DiEmph CodeArt Nova-1B Nova-6B

### v6

@@ 1 Introduction
Large language models pre-trained on source code have brought improvement in various software development domains (Chen et al., 2022a; Liu et al., 2023a; Chen et al., 2023; Le et al., 2022; Jiang et al., 2023; Xia et al., 2023). However, these LLMs are not designed for or trained with assembly corpus, not achieving their full potential on binary code analysis tasks such as binary code similarity (Wang et al., 2022; Xu et al., 2023a), malware detection (Su et al., 2024), and binary code decompilation (Tan et al., 2024; Armengol-Estapé et al., 2024; Hosseini & Dolan-Gavitt, 2022).
@@
  • We develop Nova, a generative foundation LLM with hierarchical attention and contrastive learning for assembly. Nova outperforms state-of-the-art on binary decompilation by up to 14.84 – 21.58% higher Pass@1 and Pass@10, and on binary similarity detection by up to 6.17% Recall@1.
2 Related Work
2.1 Binary Models
Machine learning models are widely used in binary program analysis tasks. However, these models are typically designed for specific tasks such as binary code similarity detection (Pei et al., 2020; Xu et al., 2023a; Wang et al., 2022; Xu et al., 2017; Ding et al., 2019), variable name prediction (Chen et al., 2022c; Xu et al., 2023b; Zhang et al., 2021; He et al., 2018; Lacomis et al., 2020), binary code type inference (Pei et al., 2021), and so on (Chen et al., 2022d; Liu et al., 2023b; Hosseini & Dolan-Gavitt, 2022).
2.2 Large Source-Code Models
LLMs demonstrate promising results on many code-related tasks, such as code generation (Chen et al., 2022a; Liu et al., 2023a; Chen et al., 2023; Le et al., 2022; Yue et al., 2021; Chen et al., 2021; Nijkamp et al., 2022; Fried et al., 2023; Rozière et al., 2023; Guo et al., 2024; Lozhkov et al., 2024; Hui et al., 2024), bug fixing (Jiang et al., 2023; Xia et al., 2023) and vulnerability fixing (Wu et al., 2023; Steenhoek et al., 2023; He & Vechev, 2023). The advances in using LLMs are attributed to the knowledge learned from massive source code and natural language text in their training datasets (Touvron et al., 2023; OpenAI, 2023). Nova is designed and trained for assembly, which has unique challenges such as low information density and diverse optimization.
2.3 Attention Mechanism
Standard self-attention is widely used in transformer architecture (Vaswani et al., 2017) to capture soft dependencies between tokens in the input. Many special attention mechanisms have been designed for better learning in various scenarios (Yang et al., 2016; Huang et al., 2024). LongCoder (Guo et al., 2023) combines window attention and global attention (attention sink (Xiao et al., 2024)) to handle long input of source code. We have shown that LongCoder’s window attention is less effective than Nova’s on assembly code. CAST (Shi et al., 2021) is a new neural architecture that splits the abstract syntax tree (AST) of source code into subtrees, encodes the subtrees, and aggregates to the final encoding. PA-former (Chai & Li, 2024) is a new neural architecture that constructs source code as pyramid input based on their AST structure and contains a pyramid attention mechanism to calculate the features in a hierarchical way. HierarchyNet (Nguyen et al., 2023) is a neural architecture that considers source code AST, data flow, and control flow graphs. Similarly, it cannot be applied to assembly code. Different from CAST, PA-former, and HierarchiyNet, Nova’s attention design is for assembly code, is more lightweight and can be plugged into any pre-trained generative LLM.
@@ 3 Approach
Figure 2 presents the overall approach of Nova. We build Nova on top of foundation models for source code (Guo et al., 2024) to utilize their source code and natural language generation ability. We first collect large assembly corpora (Section 3.1). Section 3.2 describes Nova’s hierarchical attention design. With the collected assembly corpora, we then pretrain Nova with language modeling and contrastive learning objectives (Section 3.3). Then, we fine-tune Nova on two important downstream tasks, binary code decompilation, and binary code similarity detection (Sections 3.4 and 3.5), to prove Nova’s effectiveness and benefits to the binary research domain.
@@
Table 1: Statistics (number of C and X86-64 assembly functions) of the pre-training datasets.
Origin C Functions O0 O1 O2 O3 Total
@@ 3.1 Data Collection
In this paper, we focus on X86-64 assembly functions for C programs. Yet, Nova’s approach is generalizable to other assembly languages such as ARM assembly.
We derive our X86-64 assembly functions dataset from two source code corpora: C functions in The-Stack (Li et al., 2023) and AnghaBench (da Silva et al., 2021). We compile the C programs into executables using gcc with different optimization levels (i.e., O0, O1, O2 and O3), strip the executables to remove debug information, and disassemble them into X86-64 assembly code using objdump. We treat every function as a separate data point. Table 3 shows the number of C functions in the two original datasets, and the number of X86-64 assembly functions we collected from them.
@@ 3.2 Hierarchical Self-Attention
Nova uses hierarchical self-attention that is specially designed to learn the low-information-density semantics in the long sequence of assembly code. Specifically, Nova learns the assembly code in an hierarchical way by providing a modified attention mask. Different from standard token-level attentions (Vaswani et al., 2017; Radford & Narasimhan, 2018; Radford et al., 2019; Brown et al., 2020), our hierarchical self-attention contains three categories at different levels of granularity.
@@
To sum up, the hierarchical self-attention splits assembly code semantics into three levels: intra-instruction attention captures instruction summaries, preceding-instruction attention provides context from the preceding instruction, and inter-instruction attention models long dependencies across instructions using “[INST]” tokens that contain the instruction summary. Figure 3 (c) highlights the compatibility of Nova’s hierarchical attention with standard self-attention for text and source code. Leveraging the proven performance of standard self-attention in existing LLMs, we retain the causal attention mask within and across chunks of text or source code (shown in light grey in Figure 3 (c)). Cross-attention between text, source code, and assembly is restricted to “[INST]” tokens, which encapsulate assembly instruction summaries.
@@ 3.3 Contrastive Learning
Nova is designed and implemented on decoder-only generative LLMs, and we refer the hidden states from the last transformer layer as embedding. For source code, we use the average of each token’s embedding as the source code function’s embedding. For assembly, we use the average of all the “[INST]” tokens’ embedding as the embedding of the assembly function, as each “[INST]” token is supposed to capture the semantics of that instruction by the design of our hierarchical self-attention.
Let $e^{s}_{f}$ be the embedding of function $f$ in $s$ form ($s=-1$ for source code, and $s\in[0,1,2,3]$ for O0 to O3 optimized assembly). For simplicity, let $S=[-1,0,1,2,3]$ be the domain of $s$. Functionality CL optimizes Nova’s embeddings to satisfy the constraint:
@@
The embeddings of a batch of functions, each represented in two different forms, can be optimized to satisfy these constraints. For the example in Figure 4 (b), there are two forms (source code and O0 assembly) of four functions. Once Nova encodes the batch of source code and assembly functions, we calculate the distance matrix $\{D_{ij}\}_{f_{i},f_{j}\in F}=\{d(e^{s}_{f_{i}},e^{t}_{f_{j}})\}$, and minimize the loss:
@@
This objective minimizes the distance between the embeddings of the same function, which is the diagonal in the distance matrix. In practice, this loss is computed and back-propagated for each batch of functions. Given a batch size of 64, $F$ represents a set of 64 unique functions in the batch.
Optimization CL: LLMs can be confused if being asked to directly connect a source code function to its O3-optimized assembly, due to their dramatically different syntax. Such a huge gap can be filled by learning how the source code is transformed to O0, O1, O2 and eventually to O3 assembly, as the optimization levels are ordered. Higher-level optimization applies a super-set of optimization rules compared to lower-level optimization.
Nova learns such order with the optimization CL objective, encoding differently-optimized assembly code orderly. Optimization CL optimizes Nova with the constraint: the more optimizations applied, the larger the difference between embeddings of optimized and unoptimized code. For instance, Figure 4 (c) and (d) illustrate that for the same function “cmp”, the distance between source code and assembly increases when the optimization level increases. Formally, optimization CL minimizes the following loss:
@@ 3.5 Task 2: Binary Code Similarity Detection
A widely used setting is taking a query assembly of the function $f^{q}$ that is compiled with one optimization level (denoted by $s$), and a pool of candidate assembly of $K$ (e.g., 50, 100, etc.) functions (notated by $f^{p}_{i},$ $1\leq i\leq K$) compiled with a different optimization level (denoted by $t\neq s$). There exists a unique candidate assembly coming from the same source code as the query ($\exists!1\leq i\leq K,$ $f^{p}_{i}=f^{q}$, called the positive candidate). Nova is fine-tuned to encode these binaries, so that the positive candidate has the highest similarity with the query assembly among the pool. The learning objective is as follows:
@@ 4 Experimental Setup
This section describes the setup of pre-training and fine-tuning of Nova, as well as the existing baselines we compare Nova with, and the evaluation metrics we used in the two downstream tasks. Appendix A.2 contains additional details such as training hyper-parameters, and evaluation setup.
@@ 4.1 Pre-Training
We use the C and X86-64 assembly functions collected from AnghaBench and The-Stack for pre-training. We pre-train Nova starting from DeepSeek-Coder (Guo et al., 2024), and the hierarchical attention is applied on half of the attention heads to balance between its effectiveness and the existing knowledge in the standard attention layers (Justified in Appendix A.3). Nova is pre-trained with language modeling for one epoch, followed by contrastive learning objectives for another epoch.
@@ 4.2 Fine-Tuning for Binary Code Decompilation
Baselines: Nova is compared with existing SOTA LLM4Decompile (Tan et al., 2024), Meta LLMCompiler (Cummins et al., 2024), other open-sourced general code LLMs (Rozière et al., 2023; Lozhkov et al., 2024; Guo et al., 2024; Hui et al., 2024; Mishra et al., 2024; Guo et al., 2023), and commercial LLMs GPT-3.5-Turbo and GPT-4o). LLM4Decompile trains DeepSeekCoder using the same AnghaBench corpus for binary decompilation. Meta LLMCompiler trains CodeLlama models using LLVM IRs, X86, and ARM assembly code to optimize and translate binary code.
Evaluation: We let each model sample 20 decompilations per assembly function, using the temperature of 0.2 and top_p of 0.95 (Chen et al., 2021). Except for LLM4Decompile and Nova that are fine-tuned for binary code decompilation, we provide all other baselines with three-shot examples for few-shot learning (Brown et al., 2020). The decompilations are executed with the test cases to verify the functional correctness. Finally, Pass@1 and Pass@10 (Chen et al., 2021) are reported.
@@ 4.3 Fine-Tuning for Binary Code Similarity Detection
Baselines: Nova is compared with jTrans (Wang et al., 2022), DiEmph (Xu et al., 2023a) and CodeArt (Su et al., 2024). jTrans is a Transformer (Vaswani et al., 2017) encoder trained on binaries with masked token prediction and jump target prediction tasks. DiEmph uses an instruction deemphasis technique to prevent the model from learning instruction distribution biases introduced by compilers. CodeArt proposes a regularized attention mask for encoder models to capture instructional semantics and data dependencies.
Evaluation: We randomly sample $K=$ 50, 100, 200, 500 source code functions from each project, compile them into binaries with O0 and O3 optimization flags, and disassemble them into X86-64 assemblies. BCSD techniques encode these assemblies into embeddings (Nova uses the average last-layer hidden states of all the “[INST]” tokens in an assembly as its embedding). Then each O0 assembly is used as the query to calculate their similarity with the $K$ O3 candidate assemblies. Metric Recall@1 is reported as the ratio of queries for which the candidate from the same source code has the highest similarity among all the candidates.
@@ 5.1 Binary Code Decompilation
LLM4Decompile-1.3B 15.30 8.26 9.36 8.38 10.33 21.79 15.23 16.17 13.70 16.72
Nova-1.3B 37.53 21.71 22.68 18.75 25.17 49.38 34.84 36.95 32.03 38.30
LLM4Decompile-6.7B 29.97 19.05 20.46 18.32 21.95 40.40 27.75 28.85 28.51 31.38
Nova-6.7B 48.78 30.58 30.85 27.23 34.36 57.47 47.45 43.03 39.68 46.91
@@
With the same model size, Nova-1.3B outperforms LLM4Decompile-1.3B with a 14.84% higher averaged Pass@1, and a 21.58% higher Pass@10. Nova-6.7B outperforms LLM4Decompile-6.7B with a 12.41% higher averaged Pass@1, and a 15.53% higher Pass@10. When compared with GPT-4o, an order of magnitude larger model, Nova-1.3B produces an 8.38% higher Pass@1 and 13.80% higher Pass@10. Examples of Nova’s correct decompilation are provided in Appendix A.5.
Comparison with Techniques Handling Long Input: Nova’s hierarchical attention design targets to address the low information density and long input challenge of assembly code. There are other techniques that handles long input challenges in text and source code, with Granite-3B-Code-Base-128K and LongCoder being the most related ones.
Granite trains LLM on repository-level long inputs, which is an orthogonal approach with Nova’s approach (hierarchical attention and contrastive learning). We train Granite-3B-Code-128K with Nova’s approach, and Table 3 shows that Nova’s approach brings improvement to Granite over standard fine-tuning even if it has already been trained with long code data.
@@
Granite + Nova’s Approaches 31.04 14.57 14.70 13.66 18.49 39.57 21.23 21.77 19.82 25.60
@@
Nova-1.3B (using LongCoder’s Attention) 34.59 19.07 19.72 17.34 22.68 42.19 32.37 32.86 29.04 34.12
Nova-1.3B 37.53 21.71 22.68 18.75 25.17 49.38 34.84 36.95 32.03 38.30
LongCoder combines window attention and global attention to learn long code input. We compare LongCoder’s attention design with Nova’s hierarchical attention design by replacing the hierarchical attention of Nova-1.3B with LongCoder’s attention design. Table 4 shows that Nova’s hierarchical attention is more effective in learning assembly code. Nova’s attention design considers the instruction-level local semantics and dependencies between different instructions, which fits better than fix-sized window attention to assembly code.
Ablation Study: We conduct an ablation study by comparing Nova-1.3B with the following models to show the effectiveness of contrastive learning objectives and hierarchical attention design:
  • Nova-CL-HA: Removing contrastive learning and hierarchical self-attention. This is simply training DeepSeekCoder-1.3B on the assembly corpus using language modeling. This can be viewed as our reproduction (retrain) of LLM4Decompile-1.3B using the same amount of data.
@@
Table 5: Ablation study of Nova-1.3B on HumanEval-Decompile.
@@
Benchmarks jTrans DiEmph CodeArt Nova-1.3B Nova-6.7B
@@
Benchmarks jTrans DiEmph CodeArt Nova-1.3B Nova-6.7B
@@
Benchmarks jTrans DiEmph CodeArt Nova-1.3B Nova-6.7B
@@
Benchmarks jTrans DiEmph CodeArt Nova-1.3B Nova-6.7B